# Cosmological constraints from the Minkowski functionals of the BOSS CMASS galaxy sample

### v1

@@ 1 Introduction
Current and upcoming galaxy surveys such as DESI  [^3] http://www.desi.lbl.gov, PFS  [^4] http://pfs.ipmu.jp, the Nancy Grace Roman Space Telescope  [^5] http://wfirst.gsfc.nasa.gov, Euclid  [^6] http://sci.esa.int/euclid and CSST [36, 37] [^7] http://nao.cas.cn/csst, will provide high-precision measurements of the 3D clustering of galaxies. To fully extract and exploit the wealth of valuable information provided by these cosmological observations, statistical tools beyond the standard 2-point correlation function and power spectrum are needed to probe non-Gaussian information induced by nonlinear processes. Although the 2-point statistics can be straightforwardly extended to the 3-point correlation function [38, 39] and bispectrum [40, 41, 42, 43, 44], as well as the 4-point correlation function [45, 46] and trispectrum etc.[47, 48] for the recovery of the lost non-Gaussian information. These higher-order statistics are computationally expensive, with signal-to-noise ratios that degrade rapidly at higher orders, and have a large dimensionality, so the model of their uncertainties is difficult and also computationally intensive. The challenges imposed by the n-point statistics motivate the development of various novel summary statistics and the reinforcement of previous alternative clustering methods, including but not limited to: proxies of higher order n-point statistics [49, 50, 51], the marked 2-point correlation function [52, 53, 54] or power spectrum [55, 56], the density-split clustering [57, 58, 59, 60], the nearest neighbor distributions [61, 62], the void statistics [63, 64, 65, 66, 67, 68], the counts-in-cells statistics [69, 70], the minimum spanning tree [71, 72], and the wavelet scattering transform [73, 74, 75, 76].
@@ 3.2 The halo-galaxy connection model
We have also implemented the velocity bias for both centrals $\alpha_{\rm vel,c}$ and satellites $\alpha_{\rm vel,s}$, where $\alpha_{\rm vel,c}$ describes the difference in the peculiar velocity between the central galaxy and the halo center, while $\alpha_{\rm vel,s}$ characterizes the deviation of satellite galaxies’ peculiar velocities from those of local dark matter particles. When $\alpha_{\rm vel,c}=0$ and $\alpha_{\rm vel,s}=1$, there is no velocity bias for both central and satellite galaxies. The velocity bias is a necessary extension of the HOD framework since evidence supporting the existence of velocity bias has been found in the analysis of both hydrodynamical simulations [116, 117] and observational galaxy catalogs [118, 114].
@@ 3.3 Survey systematics
Figure 2: The 3D spatial distributions of the Abacus mock produced with the pipeline described in the text for the CMASS North (first row) and South (second row) samples, the position of mock galaxies are projected along the $\hat{z}$ (first column), $\hat{y}$ (first column), $\hat{x}$ (first column) axis. The red rectangles outline the remapped simulation box, while the green squares delineate the cubic grid where the number density field is constructed from the positions of galaxies.
@@ 3.4.1 Minkowski functionals
$\sigma_{j}^{2}=\int_{0}^{\infty}\frac{k^{2}dk}{2\pi^{2}}k^{2j}P_{L}(k)e^{-(kR)^{2}/2},$ (3.6)
@@
Previous application of the MFs to the CMASS data [34, 35] focused on the amplitude of the MFs, which is determined by the ratio $\sigma_{1}/\sigma_{0}$ and only Gaussian information is extracted. In real space, the amplitude of the MFs provides a measure of the shape of the linear matter power spectrum and hence is mainly sensitive to $\omega_{\rm cdm}$ and $n_{s}$. In redshift space, the amplitude is changed by redshift space distortions and influenced by the linear growth rate and galaxy bias. We adopt a density contrast threshold parameter $\delta$ in this analysis, which is related to the one used in Eq 3.5 with $\delta=\nu\sigma$ ($\sigma$ is the variance of the smoothed density contrast field). Therefore, any factor, physical or unphysical, that changes the variance of the density field will rescale the threshold $\delta$, and the curves of the MFs as a function of $\delta$ will be expanded or compressed. This choice of threshold parameter allows us to extract more information at the cost of being prone to systematics. Most importantly, our full shape analysis of the MFs allows us to probe beyond Gaussian information in the MFs. As seen in Eq 3.5, the departure from the Gaussian prediction of the MFs is sensitive to the skewness parameters. For a non-Gaussian field like what is studied in this work, the MFs also depend on the kurtosis, and even higher-order parameters [85, 86]. The constraining power of the MFs on cosmological and HOD parameters comes from the combination of both Gaussian and non-Gaussian information embedded in the MFs.
@@
For galaxy samples with non-trial boundaries seen in figure 2, we adopt a similar algorithm to that used in [35] to obtain an unbiased estimate of the MFs. galaxies are first embedded into a uniform cubic grid of size $L_{\rm box}=2828~{}h^{-1}\rm Mpc$ and spacing $\Delta_{g}=2828/512\simeq 5.5~{}h^{-1}\rm Mpc$ for both the North and South sample. Then, we interpolate the positions of galaxies onto the cubic grid using the Cloud-in-Cell (‘CIC’)  [^13] We use the routine provided by Pylians (Python libraries for the analysis of numerical simulations [121]): https://pylians3.readthedocs.io/en/master/, other mass assignment schemes are also available, such as ‘NGP’ (nearest grid point), ‘TSC’ (triangular-shape cloud), and ’PCS’ (piecewise cubic spline) mass assignment scheme with the weight associated to each galaxy (uniform weight is used for galaxy catalogs apart from the CMASS sample and the Patchy mocks) and get a number density field $n(\boldsymbol{x})$. The survey geometry and veto masks are projected into a 3D constant field $M(\boldsymbol{x})$ with the same size and resolution as that used for the construction of the number density field, the value of $M(\boldsymbol{x})$ is set to 1 for regions inside the survey geometry but outside the veto masks, and set to 0 otherwise. For the regions with $M(\boldsymbol{x})=1$, we project the value of the angular completeness map onto them so that the 3D constant field now becomes a completeness field.
The number density field and completeness field are both smoothed with a Gaussian filter $W\left(kR_{\mathrm{G}}\right)\propto\exp\left[-k^{2}R_{\mathrm{G}}^{2}/2\right]$, we choose to work with $R_{G}=15~{}h^{-1}\rm{Mpc}$ and leave the investigation of multiple smoothing scales for future work. This smoothing scale is close to the mean galaxy separation $\bar{d}\sim 16~{}h^{-1}\rm{Mpc}$, hence this choice is aligned with previous analyses of observational data using the MFs or the genus alone141414the genus $g$ is related with the Euler characteristic $V_{3}$ by $g=1-V_{3}$ [10], the choice of $R_{G}=\bar{d}/\sqrt{2}$ [122, 6] or $R_{G}=\bar{d}$ [123, 124] is frequently used. In [125], the smoothing scale $R_{G}=\bar{d}$ is recommended for the genus topology analysis, and they found the shot noise effects diminish greatly when biased objects are used. On the other hand, the pixel size of the grid is about $5.5h^{-1}\rm{Mpc}$ in this work, which is about three times smaller than the adopted smoothing scale $R_{G}$. It is anticipated that the effect of finite pixel size [125] can be neglected for the galaxy field interpolated onto this grid [124]. To reduce the impact of complex boundaries and focus on regions with high completeness, we further discard regions where $M(\boldsymbol{x})<0.9$ after smoothing, which reduces the sample volume from $\sim(1.2~{}h^{-1}\rm{Gpc})^{3}$ to $\sim(1.0~{}h^{-1}\rm{Gpc})^{3}$. We then calculate the average number density $\bar{n}$ of galaxies and construct a density contrast field $\delta(\boldsymbol{x})=n(\boldsymbol{x})/\bar{n}-1$, within the remaining volume.
@@ 3.5 Emulators of the summary statistics
We take the galaxy mocks from cosmologies c000, c001, c002, c003, and c004 as test sets, while the remaining mocks of the first and second Latin hypercube samples described in Section 3.2 are used as the training and validation sets. That is, there are 2512, 32812, and 8197 mocks for test, training, and validation sets, respectively. The neural network emulators are trained separately for the MFs and 2PCF, with the cosmological and HOD parameters as the input layer and the concatenated four MFs or monopole and quadrupole of 2PCF as the output layer. Each element of the input parameters and the output statistics are standardized by subtracting the mean and dividing by the standard deviation. For both the emulator of MFs and 2PCF, we adopt a neural network of four hidden layers and 512 nodes per layer with a Sigmoid Linear Units activation function and a mean squared loss function. The AdamW optimization algorithm [136], an improved version of Adam [137] with the weight decay regularization correctly implemented, is utilized to optimize the weights of the neural network.
@@ 3.6 Likelihood
where $N_{\rm patchy}=2048$, $\boldsymbol{x}^{n}_{\rm patchy}$ is the measurement from the $n$th Patch mock, and $\overline{\boldsymbol{x}}_{\rm patchy}$ is the average of statistics. We find the structure of the MFs is very similar to that presented in our previous work, hence we refer the reader to [78, 81] for a detailed interpretation of the correlations between different density thresholds and between different orders of the MFs. Here we focus on the correlation between the MFs and 2PCF. There exists some correlation between the MFs and $\xi_{0}$ntermediate separations and the correlation decreases for very small and large separations. This is because the Gaussian smoothing applied to the galaxy field smears out structure much smaller than the smoothing scale $R_{G}=15~{}h^{-1}\rm Mpc$ and enables us to focus on intermediate scales around $15~{}h^{-1}\rm Mpc$ 161616we refer interested readers to figure 2 of [138] for a better understanding of the scale range probed by the MFs with a smoothing scale $R_{G}$.. On the other hand, we note that the MFs don’t correlate much with $\xi_{2}$. This is understandable: first, a considerable proportion of the velocity information is smeared out by the isotropic Gaussian smoothing process; second, the integrals 3.4 collect contributions from the infinitesimal volume or area element in an angle-independent way, this means anisotropies introduced by RSD can be buried when adding up contributions that involve the line-of-sight (LoS) direction and those that involve directions orthogonal to the LoS [91, 80]. We find the two reasons may explain why the MFs only have a weak constraining power on the velocity bias parameters, which is also discussed in Appendix C. This motivates an investigation of the Minkowski tensors, which is a straightforward extension of the MFs [139, 140], to extract anisotropies introduced by RSD [141]. We will also explore anisotropic smoothing techniques like what was used in [142] to leverage symmetries of redshift-space data and recover the lost anisotropic information in future work.
@@
where $\mathcal{P}_{\rm data}$, $\mathcal{P}_{\rm emu}$, and $\mathcal{P}_{\rm abacus}$ are the corresponding correction factor for each of the three components. We visualize the relative contributions from each of the three terms of Eq 3.17 for each order of the four MFs and each multipole moment of the 2PCF in figure 6, where $\mathcal{P}_{s}\mathbf{C}_{ii,s}$ is the diagonal element of each of the three terms on the right-hand side while $\Sigma_{s}\mathcal{P}_{s}\mathbf{C}_{ii,s}$ is the diagonal element of the covariance matrix on the left-hand side. As seen from this figure, the covariance matrix is dominated by the sample variance of the CMASS data and the emulator error. For the MFs, less than a fraction of $\sim 1/4$ error comes from the sample variance of the Abacus mocks, and the fraction can be as small as $\sim 1/5$ for $W_{3}$, the last order of the MFs. For the 2PCF, the covariance matrix is significantly dominated by the emulator error for separations smaller than $\sim 10h^{-1}\rm Mpc$.

### v2

@@ 1 Introduction
Current and upcoming galaxy surveys such as DESI  [^3] http://www.desi.lbl.gov, PFS  [^4] http://pfs.ipmu.jp, the Nancy Grace Roman Space Telescope  [^5] https://science.nasa.gov/mission/roman-space-telescope, Euclid  [^6] http://sci.esa.int/euclid and CSST [36, 37] [^7] http://nao.cas.cn/csst, will provide high-precision measurements of the 3D clustering of galaxies. To fully extract and exploit the wealth of valuable information provided by these cosmological observations, statistical tools beyond the standard 2-point correlation function and power spectrum are needed to probe non-Gaussian information induced by nonlinear processes. Although the 2-point statistics can be straightforwardly extended to the 3-point correlation function [38, 39] and bispectrum [40, 41, 42, 43, 44], as well as the 4-point correlation function [45, 46] and trispectrum etc.[47, 48] for the recovery of the lost non-Gaussian information. These higher-order statistics are computationally expensive, with signal-to-noise ratios that degrade rapidly at higher orders, and have a large dimensionality, so the model of their uncertainties is difficult and also computationally intensive. The challenges imposed by the n-point statistics motivate the development of various novel summary statistics and the reinforcement of previous alternative clustering methods, including but not limited to: proxies of higher order n-point statistics [49, 50, 51], the marked 2-point correlation function [52, 53, 54] or power spectrum [55, 56], the density-split clustering [57, 58, 59, 60], the nearest neighbor distributions [61, 62], the void statistics [63, 64, 65, 66, 67, 68], the counts-in-cells statistics [69, 70], the minimum spanning tree [71, 72], and the wavelet scattering transform [73, 74, 75, 76].
@@ 2.1 BOSS CMASS galaxy catalog
We focus on the CMASS catalog rather than both the LOWZ and CMASS catalogs because their effective redshifts differ. The effective redshift for the CMASS catalog is approximately $0.5$, while for the LOWZ catalog, it is around $0.3$. In our forward model, we aim to match the effective redshift of the CMASS data by using simulation outputs at redshift $z=0.5$, which require over 16 TB of disk space. Given the large size of the simulation dataset, downloading and analyzing the data is time-consuming. To include the analysis of the LOWZ catalog, we would need to download simulation outputs at a redshift close to $z=0.3$ and repeat the forward modeling pipeline. While we recognize the relevance of the LOWZ catalog, this work represents our first application of the MFs to observed galaxies using the simulation-based method. We plan to explore the inclusion of the LOWZ catalog in future research.
@@ 3.2 The halo-galaxy connection model
We have also implemented the velocity bias for both centrals $\alpha_{\rm vel,c}$ and satellites $\alpha_{\rm vel,s}$, where $\alpha_{\rm vel,c}$ describes the difference in the peculiar velocity between the central galaxy and the halo center, while $\alpha_{\rm vel,s}$ characterizes the deviation of satellite galaxies’ peculiar velocities from those of local dark matter particles. When $\alpha_{\rm vel,c}=0$ and $\alpha_{\rm vel,s}=1$, there is no velocity bias for both central and satellite galaxies. The velocity bias is a necessary extension of the HOD framework since evidence supporting the existence of velocity bias has been found in the analysis of both hydrodynamical simulations [116, 117] and observational galaxy catalogs [118, 114]. The velocity bias for both central and satellite galaxies must be included to explain both the Finger-of-God (FoG) effect at non-linear scales and the Kaiser effect in the linear regime, seen in the two-point correlation function of the CMASS galaxies [118]. Whether it can also faithfully explain the RSD effect, especially the FoG effect on small scales, detected by non-Gaussian statistics is an interesting question that warrants further investigation.
@@ 3.3 Survey systematics
Figure 2: The 3D spatial distribution of the Abacus mock galaxies produced with the pipeline described in the text for the CMASS North (first row) and South (second row) samples. The position of mock galaxies is projected along the $\hat{z}$ (first column), $\hat{y}$ (first column), $\hat{x}$ (first column) axis. The red rectangles outline the remapped simulation box, while the green squares delineate the cubic grid where the number density field is constructed from the positions of galaxies.
@@ 3.4.1 Minkowski functionals
$\sigma_{j}^{2}=\int_{0}^{\infty}\frac{k^{2}dk}{2\pi^{2}}k^{2j}P_{L}(k)e^{-(kR)^{2}},$ (3.6)
@@
Previous application of the MFs to the CMASS data [34, 35] focused on the amplitude of the MFs, which is determined by the ratio $\sigma_{1}/\sigma_{0}$ and only Gaussian information is extracted. In real space, the amplitude of the MFs provides a measure of the shape of the linear matter power spectrum and hence is mainly sensitive to $\omega_{\rm cdm}$ and $n_{s}$. In redshift space, the amplitude is changed by redshift space distortions and influenced by the linear growth rate and galaxy bias. We adopt a density contrast threshold parameter $\delta$ in this analysis, which is related to the one used in Eq 3.5 with $\delta=\nu\sigma$ ($\sigma$ is the standard deviation of the smoothed density contrast field). Therefore, any factor, physical or unphysical, that changes the variance of the density field will rescale the threshold $\delta$, and the curves of the MFs as a function of $\delta$ will be expanded or compressed. This choice of threshold parameter allows us to extract more information at the cost of being prone to systematics. Most importantly, our full shape analysis of the MFs allows us to probe beyond Gaussian information in the MFs. As seen in Eq 3.5, the departure from the Gaussian prediction of the MFs is sensitive to the skewness parameters. For a non-Gaussian field like what is studied in this work, the MFs also depend on the kurtosis, and even higher-order parameters [85, 86]. The constraining power of the MFs on cosmological and HOD parameters comes from the combination of both Gaussian and non-Gaussian information embedded in the MFs.
@@
For galaxy samples with non-trial boundaries seen in figure 2, we adopt a similar algorithm to that used in [35] to obtain an unbiased estimate of the MFs. The galaxies are first embedded into a uniform cubic grid of size $L_{\rm box}=2828~{}h^{-1}\rm Mpc$ and spacing $\Delta_{g}=2828/512\simeq 5.5~{}h^{-1}\rm Mpc$ for both the North and South sample. Then, we interpolate the positions of galaxies onto the cubic grid using the Cloud-in-Cell (‘CIC’)  [^13] We use the routine provided by Pylians (Python libraries for the analysis of numerical simulations [121]): https://pylians3.readthedocs.io/en/master/, other mass assignment schemes are also available, such as ‘NGP’ (nearest grid point), ‘TSC’ (triangular-shape cloud), and ’PCS’ (piecewise cubic spline) mass assignment scheme with the weight associated to each galaxy (uniform weight is used for galaxy catalogs apart from the CMASS sample and the Patchy mocks) and get a number density field $n(\boldsymbol{x})$. The survey geometry and veto masks are projected into a 3D constant field $M(\boldsymbol{x})$ with the same size and resolution as that used for the construction of the number density field, the value of $M(\boldsymbol{x})$ is set to 1 for regions inside the survey geometry but outside the veto masks, and set to 0 otherwise. For the regions with $M(\boldsymbol{x})=1$, we project the value of the angular completeness map onto them so that the 3D constant field now becomes a completeness field.
The number density field and completeness field are both smoothed with a Gaussian filter $W\left(kR_{\mathrm{G}}\right)\propto\exp\left[-k^{2}R_{\mathrm{G}}^{2}/2\right]$, we choose to work with $R_{G}=15~{}h^{-1}\rm{Mpc}$ and leave the investigation of multiple smoothing scales for future work. This smoothing scale is close to the mean galaxy separation $\bar{d}\sim 16~{}h^{-1}\rm{Mpc}$, hence this choice is aligned with previous analyses of observational data using the MFs or the genus alone141414the genus $g$ is related with the Euler characteristic $V_{3}$ by $g=1-V_{3}$ [10], the choice of $R_{G}=\bar{d}/\sqrt{2}$ [122, 6] or $R_{G}=\bar{d}$ [16, 123] is frequently used. We can in principle explore down to the resolution scale of the simulations with our simulation-based method. However, it is difficult to faithfully forward model all observational systematics on very small scales. As we mentioned in Sec 3.3, we rely on the weights suggested in [96] to account for imaging systematics, redshift failures, and fiber collisions. Our main concern is that the weight used to correct for the fiber collision effect in the CMASS galaxies may not be effective on small scales [124]. We believe a comprehensive study of the impact of fiber collisions on higher-order statistics like the MFs, and how to properly correct or forward model this effect, is essential before going to smaller scales. However, this is beyond the scope of the current paper and is an area we are actively investigating. On the other hand, we find using a smaller scale doesn’t necessarily guarantee tighter parameter constraints for the CMASS galaxy sample due to the low number density ($n(z)=2.4\times 10^{-4}~{}h^{3}\rm{Mpc}^{-3}$) adopted in this analysis. The larger shot noise on smaller scales can make the curves of the MFs more noisy, which leads to a lower accuracy of the emulator. The DESI LRG sample [125] has roughly twice this number density, allowing for the extraction of more non-Gaussian information on smaller scales.
To reduce the impact of complex boundaries and focus on regions with high completeness, we further discard regions where $M(\boldsymbol{x})<0.9$ after smoothing, which reduces the sample volume from $\sim(1.2~{}h^{-1}\rm{Gpc})^{3}$ to $\sim(1.0~{}h^{-1}\rm{Gpc})^{3}$. We then calculate the average number density $\bar{n}$ of galaxies and construct a density contrast field $\delta(\boldsymbol{x})=n(\boldsymbol{x})/\bar{n}-1$, within the remaining volume.
@@ 3.5 Emulators of the summary statistics
We take the galaxy mocks from cosmologies c000, c001, c002, c003, and c004 as test sets, while the remaining mocks of the first and second Latin hypercube samples described in Section 3.2 are used as the training and validation sets. That is, there are 2512, 32812, and 8197 mocks for test, training, and validation sets, respectively. The neural network emulators are trained separately for the MFs and 2PCF, with the cosmological and HOD parameters as the input layer and the concatenated four MFs or monopole and quadrupole of 2PCF as the output layer. Due to the significant variations in the amplitude of the MFs across different orders and in the 2PCF for different separations, it is beneficial to first scale them to a comparable range. This ensures that each data point contributes equally to the loss function, which measures the root mean squared error between each element of the input statistics and the emulator prediction. We standardize both the input parameters and output statistics by subtracting the mean and dividing by the standard deviation. Additionally, we have experimented with scaling the statistics to a range between 0 and 1 and/or using a mean squared error loss function weighted by the variance of the statistics. We find that these alternative approaches do not significantly affect our results.
For both the emulator of MFs and 2PCF, we adopt a neural network of four hidden layers and 512 nodes per layer with a Sigmoid Linear Units activation function. The AdamW optimization algorithm [136], an improved version of Adam [137] with properly implemented weight decay regularization, is employed to optimize the network’s weights. The neural network is trained with a batch size of 64 and a learning rate of $10^{-3}$. To address learning stagnation, we reduce the learning rate by a factor of 2 every 30 epochs if the validation loss does not improve, continuing this adjustment until the learning rate reaches its minimum value of $10^{-6}$. The values of these hyperparameters of both the neural network and its training process are chosen based on insights from our previous work [60, 75]. We also explored alternative hyperparameter choices for the neural network and its training process using Optuna  [^16] https://github.com/optuna/optuna to minimize the validation loss. After 200 trials with Optuna, varying the learning rate, number of layers, number of hidden units per layer, dropout rate, and weight decay, we found no improvement over the emulator trained with our default settings.
@@ 3.6 Likelihood
where $N_{\rm patchy}=2048$, $\boldsymbol{x}^{n}_{\rm patchy}$ is the measurement from the $n$th Patch mock, and $\overline{\boldsymbol{x}}_{\rm patchy}$ is the average of statistics. We find the structure of the MFs is very similar to that presented in our previous work, hence we refer the reader to [78, 81] for a detailed interpretation of the correlations between different density thresholds and between different orders of the MFs. Here we focus on the correlation between the MFs and 2PCF. There exists some correlation between the MFs and $\xi_{0}$ of intermediate separations and the correlation decreases for very small and large separations. This is because the Gaussian smoothing applied to the galaxy field smears out structure much smaller than the smoothing scale $R_{G}=15~{}h^{-1}\rm Mpc$ and enables us to focus on intermediate scales around $15~{}h^{-1}\rm Mpc$ 171717we refer interested readers to figure 2 of [32] for a better understanding of the scale range probed by the MFs with a smoothing scale $R_{G}$.. On the other hand, we note that the MFs don’t correlate much with $\xi_{2}$. This is understandable: first, a considerable proportion of the velocity information is smeared out by the isotropic Gaussian smoothing process; second, the integrals 3.4 collect contributions from the infinitesimal volume or area element in an angle-independent way, this means anisotropies introduced by RSD can be buried when adding up contributions that involve the line-of-sight (LoS) direction and those that involve directions orthogonal to the LoS [91, 80]. We find the two reasons may explain why the MFs only have a weak constraining power on the velocity bias parameters, which is also discussed in Appendix C. This motivates an investigation of the Minkowski tensors, which is a straightforward extension of the MFs [138, 139], to extract anisotropies introduced by RSD [140]. We will also explore anisotropic smoothing techniques like what was used in [141] to leverage symmetries of redshift-space data and recover the lost anisotropic information in future work.
@@
where $\mathcal{P}_{\rm data}$, $\mathcal{P}_{\rm emu}$, and $\mathcal{P}_{\rm abacus}$ are the corresponding correction factors for each of the three components. Their values are 1.46, 1.34, and 1.55, respectively. We visualize the relative contributions from each of the three terms of Eq 3.17 for each order of the four MFs and each multipole moment of the 2PCF in figure 6, where $\mathcal{P}_{s}\mathbf{C}_{ii,s}$ is the diagonal element of each of the three terms on the right-hand side while $\Sigma_{s}\mathcal{P}_{s}\mathbf{C}_{ii,s}$ is the diagonal element of the covariance matrix on the left-hand side. As seen from this figure, the covariance matrix is dominated by the sample variance of the CMASS data and the emulator error. For the MFs, less than a fraction of $\sim 1/4$ error comes from the sample variance of the Abacus mocks, and the fraction can be as small as $\sim 1/5$ for $W_{3}$, the last order of the MFs. For the 2PCF, the covariance matrix is significantly dominated by the emulator error for separations smaller than $\sim 10h^{-1}\rm Mpc$.